I am having so much trouble with this problem, and I need help. I have tried and tried, but I am not having any luck! Any help would be greatly appreciated! Thank you!

Solve, and please show work so I may learn how to do these types of problems.
If Brenda can clean the house in 2 hours and Kenny can clean the house in 3.5 hours, how long will it take them to clean the house if they work together?

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

THE SOLUTION TO THESE TYPE OF PEOBLEMS IS THE PRODUCT OF THE TWO TIMES DIVIDED BY THE SOME OF THE TWO TIMES OR (3.5*2)/(3.5+2) OR 7/5.5 OR 1.27 HOURS

Answer by stanbon(75887)

If Brenda can clean the house in 2 hours and Kenny can clean the house in 3.5 hours, how long will it take them to clean the house if they work together?
Brenda Data:
time = 2 hrs; rate= (1/2)job/hr.
Kenny Data:
time = (7/2) hrs; rate = (2/7)job/hr.
Together Data
time = x hrs; rate = (1/x)job/hr.
EQUATION:
rate + rate = rate
1/2 + 2/7 = 1/x
Multiply both sides by 14x to get:
7x + 4x = 14
11x = 14
x= 14/11 = 1.273 hrs = 1 hr 17 minutes
Cheers,
Stan H.